help with the top part of the question
Answer #1

1 mole of O2 is formed by passage of four Faraday (4F)of electricity.

The quantity of charge when 81.8 A ampere is passed for 83.7 minutes = 81.8*83.7*60 =410799.6 Coulomb's of charge. Therefore, the number of Faradays of charge 410799.6/96500 = 4.26 F.

The number of moles of O2 gas produced by the passage of 4.26 F is 4.26*1/4 = 1.06 moles.

At STP conditions, 1 mole of O2 gas occupies 22.4 liters of volume. Therefore, the volume occupied by 1.06 moles of O2 gas 1.06*22.4 = 23.84 liters.